OpenStudy (anonymous):

Consider the function f(x)=x^2+5x+5 and the point A(4,5). f′(x)=_______ Find a formula for the slope of a line passing through point A, and an arbitrary point on the function. Your answer should be a formula in terms of x. slope=_________ Using your answers from the two previous questions, find the slope(s) of the line(s) through point A that is (are) tangent to f(x). If there is more than one line you can separate your answers with a comma (ex. 5,2 ). slope(s)=_____________

OpenStudy (anonymous):

I know that f'(x) = 2x+5. But I dont know where to start with the second and third part. Any help would be greatly appreciated!!

OpenStudy (anonymous):

u r right f'(x) = 2x+5.

OpenStudy (anonymous):

(5-(x^2+5x+5 ))/(4-x) =(-x^2-5x)/4-x =(x^2+5x)/x-4

OpenStudy (anonymous):

you could substitute x as 4 in 2x+5, so that you'll get the f'(x). which is equivalent to slope

OpenStudy (anonymous):

@darkmare got it that's the answer for second question

OpenStudy (anonymous):

how did you know what to do to find the equation for the slope?

OpenStudy (anonymous):

where did the x^2+5x_5 come from?

OpenStudy (anonymous):

simple one point is A(4,5) arbitary point is (x,y) i.e (x,x^2+5x+5 ) find the slope of these two points

OpenStudy (anonymous):

oooo ok

OpenStudy (anonymous):

so how would I figure out the last part?

OpenStudy (anonymous):

is the answer for last part is 13

OpenStudy (anonymous):

no i know its not 13

OpenStudy (anonymous):

k just wait i will find it

OpenStudy (anonymous):

for #3, answers are 25,1

OpenStudy (anonymous):

I just figured that out :) Thanks for your help guys!! :)

OpenStudy (anonymous):

right